OpenStudy (anonymous):

what are the roots of the polynomial x^2-3x+5?

OpenStudy (anonymous):

No real roots

OpenStudy (anonymous):

what are the roots thougth? is one of them -3+sqrt11/2?

OpenStudy (anonymous):

[3 + i(root 11)]/2

OpenStudy (amistre64):

(x -3/2)^2 = -5 + 9/4 x -3/2 = +-sqrt(-11)/2 x = \(\frac{3}{2} \pm i {\sqrt{11}\over2}\) maybe?

OpenStudy (anonymous):

replace the + wirh a - for the second root

OpenStudy (anonymous):

see gblover, its nice to see ur attempting,so here's a new concept... actually u'll get, (-3+-sqrt(-11))/2 as the roots...but u seems to hav ignored the -ve sign in sqrt(-11),which is wrong...as amistre did u need to follow that. there i is called the imaginary number coefficient .